OpenStudy (anonymous):

Which expression is equivalent to 8a + 20? 4(2a + 20) 4(2a + 5) 4(4a + 16) 4(4a + 5 )

jimthompson5910 (jim_thompson5910):

Pick an answer choice at random. Say choice D Distribute the 4 through: 4(4a + 5 ) = 4*4a + 4*5 = 16a + 20 so we can rule out choice D because we want 8a+20 and not 16a+20

OpenStudy (arindameducationusc):

just facotr it. you will get 4(2a+5)

OpenStudy (arindameducationusc):

you there?